Astronomy Report – December 15th

Astronomy Report

Name:  Thomas Horn    Crew: 184
Date: 7/15/17

Sky Conditions: Clear

Wind Conditions: Light

Observation Start Time: 2:00 pm

Observation End Time: 2:30 pm

Summary: No observing, shutdown Observatory and returned Laptop to HAB

Objects Viewed: None

Problems Encountered: None

Sol Summary – December 15th

MDRS Sol Summary Report for Sol 14

Summary Title: Signing off!

Mission Status:

It’s been a please working with all of you, and thanks for your help.

It was a fitting send off to our team that our day started with a
generator failure. Upon our morning rounds after waking up it was
noticed the Generator had failed with an ‘Underspeed’ Fault. We were
happy to see that SOC was holding steady at 100% (thank you solar
arrays!) but this was a worrying sign, especially as we didn’t want to
hand over a crippled ship to the oncoming crew.

Investigation into the generator showed that the loose connections
that were repaired previously were still holding strong, so there was
no immediate cause of the generator fault. After attempting to
restart the generator a low battery voltage was observed (reading
7.6volts, supposed to be 9.6 volts). It appears that after the
generator failed it drained the battery down displaying the fault
message on its status plate. After a jumpstart from our friendly
neighbors we were able to get the generator back up and running. This
issue is likely to re-occur as we are still not sure of the original
cause. Our next step will be to Remove and Replace the generator
batteries, fingers crossed this fixes the problem. Best of luck to
future crews in continuing to troubleshoot this issue.

Besides that morning excitement the day was one of cleaning and
packing, with a short break in the afternoon to do a short EVA around
the HAB for some last minute filming to get the shots we’ve missed so
far this mission. We hope to leave this place better than we found
it.

This report comes to you with the assistance of Mission Support as our
Internet allowance has run out. Sometime today the same laptop that
killed our bandwidth before had WiFi turned On, after which it sucked
our bandwidth dry again. Still don’t know why this particular laptop
uses bandwidth at such a prodigious rate, but we’ve turned it off
again and comms should be returned to normal.

Hope to work with you all again in the future.

Ad Astra Per Ardua,

Crew 184.

Sol Activity Summary:

1. Filming EVA

2. Cleaning / Packing

4. Reports and general HAB maintenance / checks

6. Generator Failure and Jumpstart

Look Ahead Plan:

1. Crew Handover Training

Anomalies in work:

1. HAB Leak from under EVA door continuing (per Mission Control
leak is expected and has been present for months, presumed to be from
Kitchen piping)

2. Generator Underspeed Fault

Weather: Cold, but temps are above freezing in mid-day

Crew Physical Status: All crew in good physical health.

EVA: None today.

Reports to be filed:

Greenhab

Operations

Journalist

Astronomy

Sol Summary

Mission Summary

Support Requested:

1. None

Operations Report – December 15th

Crew 184 Operations Report 12/15/2017

SOL: SOL 14

Name of person filing report: JHunt

Non-nominal systems: Generator

Notes on non-nominal systems:  Generator was found turned off at crew
wake. Status light was red and Fault Message read “Underspeed Fault”.
Last time generator was confirmed ON was around 7:00am. It is unknown
what caused the sudden power-down, but the solution was a jumpstart
from the Shuttle battery (Shannon’s Truck). That happened around
3:00pm and generator is running nominally at this time.

Generator (hours run): Generator turned off at Unknown time (failure).

Generator turned on at 3:00pm.

Solar— SOC @ 12:00pm : 100%

SOC @ 3:00 pm: 100%

Diesel – 85%

Propane – 64 percent volume

Ethanol Free Gasoline (5 Gallon containers for ATV) – 5 Gallons (Full)

Water (trailer) – 400 gallons

Water (static) – 250 gallons

Trailer to Static Pump used – No

Water (loft) – Static to Loft Pump used – Yes

Water Meter: 46 gallons

Toilet tank emptied: No

ATV’s Used: None

Oil Added? No (Checked oil. All was nominal)

ATV Fuel Used:

# Hours the ATVs were Used today: 00:00 hours

Notes on ATVs: ATVs were nominal except, of course, the Yamaha 350
with the bad battery. Battery was actually removed from ATV today and
disposed of.

Deimos rover used: No

Hours: 00:00 hours

Beginning charge:

Ending charge:

Currently charging:Yes

Sojourner rover used:  ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used:  No

Hours: 00:00 Hours

Beginning charge:

Ending charge:

Currently charging: Yes

Opportunity rover used:  No

Hours: 00:00 Hours

Beginning charge:

Ending charge:

Currently charging: Yes

Curiosity rover used:  Not Used

Hours:

Beginning charge:

Ending charge:

Currently charging: Yes

HabCar used and why, where?  No

Summary of internet: Internet is out. Unsure what caused the absolute drop.

Summary of suits and radios: All Nominal

Summary of Hab operations: All nominal

Summary of GreenHab operations: Greenhab report is nominal and cleaned!

Summary of ScienceDome operations: All nominal

Summary of RAMM operations: Not Operational

Summary of health and safety issues: Crew is Healthy

Questions, concerns and requests to Mission Support:

GreenHab Report – December 15th

GreenHab Report

Trisha Randazzo

December 15, 2017

Environmental control: (Choose which is appropriate and explain further if needed)

Ambient (no heating or/cooling)

Ambient with window/door open

·         Outside temperature during work hours:  10 C

·         At 12:00 inside temperature was 43 C. Opened door to cool.

·         Keep door open for most of the day. Closed around 16:30, inside temperature was 20 C.

·         Inside temperature high 43 low 18

·         Inside humidity: 16%

Heating

·         Functioning nominally

·         Turns on at 16 C stops at 20 C

Cooling

·         Not available

Both heating and cooling

·         Nominal

Shade cloth on/off

·         On

Average temperature: (N/A until new sensor is delivered)

·         N/A

Changes to crops:  Note all emergence of seedlings, death of seedlings, etc

·         Three of the bean plants are still drooping

·         One of the bean plants is flowering! A nice purple flower.

·         Other beans are starting to appear on 4 bean plants!

·         No sprouts yet from the newly planted herb seeds. Expect some by next Monday!

Daily water usage for crops:

·         Used more water to accommodate the drier planter box.

Time(s) of watering for crops:

·         15:00

Morning research observations: (Currently not operational)

·         N/A

Changes to research plants:

·         N/A

Daily watering and amount of water used:

·         3.0 gallons

Aquaponics:  (Currently not operational)

·         N/A

Narrative:  Any other information you want to share

·         After yesterday’s report sent, Mission Support delivered  more seeds and plant food. Thank you mission support!

Support/supplies needed:

·         N/A